# Can MLLMs Guide Weakly-Supervised Temporal Action Localization Tasks?

### v1

@@ Abstract
Introduction
Temporal action localization (TAL) (Paul, Roy, and Roy-Chowdhury 2018; Narayan et al. 2021; Lee et al. 2021; Alayrac et al. 2022; Dai et al. ) aims to localize action instances of interest from untrimmed videos. Although current fully-supervised TAL methods achieve significant localization results, fully-supervised methods require expensive, time-consuming, and labor-intensive frame-level annotations, and weakly-supervised temporal action localization (WTAL) methods (Paul, Roy, and Roy-Chowdhury 2018; Narayan et al. 2021; Lee et al. 2021), which require only video-level labels, have received much attention in order to alleviate the cost of annotation.
Existing WTAL methods typically train a classifier(Paul, Roy, and Roy-Chowdhury 2018; Narayan et al. 2021; Lee et al. 2021) using video-level labels, generating a set of probability predictions. This process results in temporal class activation maps (T-CAM). Despite notable performance improvements, current methods still face two challenges: incomplete and over-complete localization. As illustrated in Fig.1, some sub-actions with low discriminability may be ignored, leading to incomplete localization, while certain background segments that contribute to classification might be misclassified as foreground actions, causing over-complete localization.
@@ Introduction
To overcome these challenges, we turn our perspective to existing multimodal large language models(MLLMs)(Alayrac et al. 2022; Dai et al. ). The state-of-the-art MLLM employs a seamless amalgamation of a video foundation model and a large language model, culminating in the development of a sophisticated system for enhanced video comprehension. This integrated framework transcends the limitations inherent in traditional predefined visual tasks. Although MLLM exhibits powerful video understanding capabilities, the resource-intensive nature of running it poses high memory and computational costs. Therefore, traditional models remain essential for video understanding tasks. Is it possible to propose a learning paradigm that leverages MLLMs to enhance existing traditional WTAL methods? Inspired by this, we introduce a paradigm for MLLM-enhanced WTAL tasks, known as MLLM4WTAL.
@@
Related Work
Temporal Action Localizationl requires the simultaneous prediction of action start and end times along with the corresponding category. In fully-supervised temporal action localization training, frame-level labeling is commonly employed. Existing methods fall into two main categories: top-down and bottom-up. Top-down methods (Chao et al. 2018; Gao et al. 2017) transform action localization into a time-dimensional detection task, inspired by image object detection. These methods generate action proposals, followed by classification and regression on temporal boundaries. They can draw on state-of-the-art techniques for Object Detection. On the other hand, bottom-up methods (Xu et al. 2020), (Bai et al. 2020) generate frame-level predictions, followed by action localization using cleverly designed post-processing techniques. However, this fully supervised approach relies heavily on frame-by-frame annotation, which is both expensive and impractical when dealing with long videos.
Weakly-Supervised Temporal Action Localization has garnered increasing attention within the research community, particularly due to its ability to learn from video-level category labels during training, thus circumventing the need for precise temporal annotations(Paul, Roy, and Roy-Chowdhury 2018; Narayan et al. 2021; Lee et al. 2021). The initial approach was UntrimmedNet (Wang et al. 2017), which first addressed this problem by employing a Multiple Instance Learning (MIL) framework that selects foreground segments and combines them into action segments. Subsequent improvement methods, such as STPN (Nguyen et al. 2018), enhanced UntrimmedNet by introducing a sparse loss to make the selected segments more sparse. CoLA (Zhang et al. 2021) utilizes contrastive learning to distinguish between foreground and background segments. In addition, UGCT (Yang et al. 2021) proposed an online pseudo-label generation mechanism with uncertainty-aware learning capability for pseudo-label supervision of attention weights.
Multimodal Large Language Models have emerged as a powerful paradigm for tackling a wide range of tasks spanning multiple modalities. LLM (Brown et al. 2020; Chiang et al. 2023; OpenAI 2023; Touvron et al. 2023) has recently excelled in Natural Language Processing (NLP) tasks. To integrate with other modalities, researchers have worked to construct MLLM (Li et al. 2023a, c; Zhu et al. 2023). Flamingo (Alayrac et al. 2022) integrates pre-trained visual and linguistic models to achieve state-of-the-art performance using a small number of samples. BLIP-2 (Li et al. 2023c) proposes a generic and efficient pre-training strategy that utilizes frozen pre-trained image encoders and large linguistic models for visual language pre-training. MiniGPT4 (Zhu et al. 2023) implements the system by aligning frozen visual coders and LLMs, Vicuna (Chiang et al. 2023), using only one projection layer. Otter (Li et al. 2023a) demonstrates improved instruction-following and context-learning capabilities. In the domain of video, ChatVideo (Wang et al. 2023) implements user interaction with LLMs using the tracklet as the fundamental video unit. VideoChat (Li et al. 2023d) integrates the video foundation models and LLMs through a learnable neural interface and performs well in spatio-temporal reasoning, event localization, and causal relationship inference. As for Video-LLaMA (Zhang, Li, and Bing 2023), it utilizes the pre-trained models ImageBind (Girdhar et al. 2023) and LLaMA (Touvron et al. 2023) on top of BLIP-2, which further facilitates cross-modal training for video.
@@
Methods
@@
Overview.The overall framework of proposed MLLM4WTAL is illustrated in Fig.2. It leverages the semantic priors of MLLM to enhance WTAL from two aspects: Key Semantic Matching (KSM) and Complete Semantic Reconstruction (CSR). The framework employs a dual priors interactive enhancement strategy to achieve interactive enhancement of these two semantic priors. For section Key Semantic Matching, MLLM first takes the raw video input to obtain a temporal action key description. Subsequently, the key semantic query prompt is constructed based on the prompt template and input to the text encoder to get the key semantic query. Then RGB and flow video features $X_{rgb}$ and $X_{flow}$ are fed into the video embedding module to generate video feature embedding. In the key semantic matching module, KSM matches the key semantic query with all video segments, obtaining a contrastive query response that activates the video segment most matching the key semantic. Additionally, attention weights are generated for each video segment to further suppress the response of background segments to the key semantic activation. For section Complete Semantic Reconstruction, first MLLM inputs the raw video to get a complete description of the temporal action, then masks the sentence’s key action words. Next, the attention mechanism is designed to collect all the complete fragments associated with the masked action words and predict the masked action words by a linguistic reconstructor. Finally, in section Dual Prior Interactive Distillation, a dual priors interactive enhancement strategy is employed to achieve a strong collaboration between KSM and CSR, aiming to obtain more accurate and complete localization results.
Key Semantic Matching
In this section, we will introduce key semantic matching (KSM) to comprehensively activate key temporal action intervals. Specifically, KSM consists of a key semantic generator, video embedding module, text embedding module, and semantic matching module. Key Semantic Generator. In the key semantic generator module, the first step is to construct key semantic prompts for MLLM, guiding the MLLM to generate key descriptions for temporal actions. It is worth noting that when building the prompts, we have prior knowledge of the action categories present in the video. We introduce this action category before our prompt template. The prompt template $P_{\text{key}}$ format is as follows:
@@
Video Embedding Module. Similar to other WTAL models, the video embedding module consists of two 1D convolutions followed by ReLU and Dropout layers. We adopt the same strategy as in (Bai et al. 2020) to fuse the RGB and flow features to obtain the input $\mathrm{F}\in\mathrm{R}^{\mathrm{T}\times 2048}$ of the video embedding module.
@@
The corresponding video feature embedding $\mathrm{F}_{\mathrm{e}}\in\mathrm{R}^{\mathrm{T}\times 2048}$ can then be obtained by $F_{e}=\Phi_{\text{emb}}(F)$, where $\Phi_{\text{emb}}$ represents the video embedding module. In addition, following previous works (Hong et al. 2021; Islam, Long, and Radke 2021), an attention mechanism is used to generate attention weight $\mathcal{A}_{KSM}\in\mathrm{R}^{\mathrm{T}\times 1}$ for each video clip.
@@
where $T_{\text{start}}$ represents the [START] token that was randomly initialized, $T_{\text{context}}$ is the learnable textual context of length $N_{context}$, and $T_{\text{key}}$ is the key informative textual feature embedded with GloVe (OpenAI 2023). Additionally, the $(C+1)th$ additional context class embedding is initialized with 0. Subsequently, a transformer encoder, denoted as $\Phi_{\text{trans}}$, functions as the text embedding module to generate text queries. Importantly, it should be noted that the key semantic query $F_{\text{query}}$ can be obtained through $F_{\text{query}}=\Phi_{\text{trans }}\left(T_{\text{query}}\right)$, where $F_{\text{query}}\in\mathrm{R}^{(\mathrm{C+1})\times 2048}$.
Semantic Matching Module The semantic matching module aims to activate video segment features associated with key semantic queries. Specifically, we perform inner product operation on video embedding feature $X_{e}$ and text embedding feature $X_{q}$ to generate segment-level video-text similarity matrix $M\in R^{T\times(C+1)}$. In addition to this, following (Hong et al. 2021; Islam, Long, and Radke 2021), we also employ the attention weight $\mathcal{A}_{KSM}$ to suppress the background segment responses. The segment-level matching result $\hat{M}\in R^{T\times(C+1)}$ for background suppression can be obtained by $\hat{M}=\mathcal{A}_{KSM}*M$, where ${}^{\prime}*^{\prime}$ denotes the element-wise multiplication. Following (Paul, Roy, and Roy-Chowdhury 2018; Min and Corso 2020), we employ top-k multi-instance learning to compute the matching loss. Specifically, we compute the average of top-k similarities on the time dimension of the response of a particular text query category as the video-level video-text similarity. For the j-th action category, the video-level similarities $\mathrm{S}_{\mathrm{j}}$ and $\hat{S}_{\mathrm{j}}$ are generated from $M$ and $\hat{M}$, respectively.
@@
Complete Semantic Reconstruction
@@
To explore the complete temporal intervals of video that are semantically related to text, CSR includes an attention mechanism similar to Key Semantic Matching. The attention weights $\mathcal{A}_{CSR}$ of the CSR module can be obtained through the following equation:
@@
Text Embedding Module. We mask the key action words of the video-complete description and embed the masked sentences with GloVe(Pennington, Socher, and Manning 2014) and a fully connected layer to obtain the complete semantic embedding $\mathrm{F}_{\mathrm{c}}\in\mathrm{R}^{\mathrm{m}\times 512}$, where m is the length of the sentence.
Transformer Reconstructor. The Transformer reconstructor is used to reconstruct masked description sentences in the Complete Semantic Reconstruction module. First, Following (Zhai et al. 2020), we randomly mask 1/3 of the words in the sentence as alternative description sentences, which may lead to a high probability of masking action labeling text. Then, the transformer’s encoder is utilized to obtain foreground video features $\mathrm{F}_{fg}\in\mathrm{R}^{\mathrm{T}\times 512}$
@@
$\mathbf{P}\left(w_{i}\mid\mathbf{F}_{\text{complete}},\mathbf{F}_{c[0:i-1]}\right)=\operatorname{sigmoid}\left(\Phi_{\mathrm{FC}}(\mathbf{H})\right)$ (16)
Where $\Phi_{\mathrm{FC}}$ is the fully connected layer and $N_{v}$ is the word size.
@@
Dual Prior Interactive Distillation
Following PivoTAL(Nayeem et al. ), we utilized a prior-guided localization head. The loss function for the localization head is as follows:
@@
Full SSN(Zhao et al. 2017) ICCV2017 66.0 59.4 51.9 41 29.8 - - 55.7 - -
P-GCN(Zeng et al. 2019) CVPR2019 69.5 67.8 63.6 57.8 49.1 - - 61.6 - -
G-TAD(Long et al. 2019) CVPR2020 - - 54.5 47.6 40.2 30.8 23.4 - 39.3 -
ActionFormer(Zhang, Wu, and Li 2022) ECCV2022 - - 82.1 77.8 71.0 59.4 43.9 - 66.8 -
Weak CoLA(Zhang et al. 2021) CVPR2021 66.2 59.5 51.5 41.9 32.2 22.0 13.1 50.3 32.1 40.9
DCC(Li et al. 2022) CVPR2022 69.0 63.8 55.9 45.9 35.7 24.3 13.7 54.1 35.1 44.0
Li et al.(Li et al. 2023b) CVPR2023 71.1 65.0 56.2 47.8 39.3 27.5 15.2 55.9 37.2 46.0
Wang et al.(Wang, Li, and Wang ) CVPR2023 73.0 68.2 60.0 47.9 37.1 24.4 12.7 57.2 36.4 46.2
Ren et al.(Ren et al. 2023) CVPR2023 71.8 67.5 58.9 49.0 40.0 27.1 15.1 57.4 38.0 47.1
STCL-Net(Fu, Gao, and Xu ) TPAMI2023 72.7 67.1 58.2 49.7 41.8 28.7 16.0 57.9 38.9 47.7
Ju et al.(Liu et al. b) CVPR2023 73.5 68.8 61.5 53.8 42.0 29.4 16.8 59.9 40.7 49.4
PivoTAL(Nayeem et al. ) CVPR2023 74.1 69.6 61.7 52.1 42.8 30.6 16.7 60.1 40.8 49.7
AHLM(AHL 2023) ICCV2023 75.1 68.9 60.2 48.9 38.3 26.8 14.7 58.3 37.8 47.6
CASE(Liu et al. a) ICCV2023 72.3 67.1 59.2 49.4 37.7 24.2 13.7 57.1 36.8 46.2
ISSF(Yun et al. 2023) AAAI2024 72.4 66.9 58.4 49.7 41.8 25.5 12.8 57.8 37.6 46.8
[…]
@@
Zhou et al.(Zhou et al. 2023) CVPR2023 74.0 69.4 60.7 51.8 42.7 26.2 13.1 59.7 38.9 48.3
MLLM4WTAL (Zhou et al.) - 74.3 69.8 61.8 52.3 43.0 30.8 16.6 61.7 40.9 49.8
@@
CoLA(Zhang et al. 2021) CVPR2021 42.7 25.7 5.8 26.1
ACGNet(Lee et al. 2021) AAAI2022 41.8 26.0 5.9 26.1
DGCNN(Shi et al. 2022) MM2022 42.0 25.8 6.0 26.2
P-MIL(Ren et al. 2023) CVPR2023 44.2 26.1 5.3 26.5
ECM(Zhao et al. 2022) TPAMI2023 41.0 24.9 6.5 25.5
Ren et al.(Ren et al. 2023) CVPR2023 44.2 26.1 5.3 25.5
STCL-Net(Fu, Gao, and Xu ) TPAMI2023 44.0 26.1 5.3 26.6
Ju et al.(Liu et al. b) CVPR2023 48.3 29.3 6.1 29.6
CASE(Liu et al. a) ICCV2023 43.8 27.2 6.7 27.9
[…]
@@
Table 3: Evaluation of each component of our method
@@
THUMOS14(Idrees et al. 2017) contains 413 untrimmed videos covering 20 different categories of actions. On average, each video contains about 15 unique instances of actions. We trained using 200 validation videos and evaluated on 213 test videos. Despite the small size of the dataset, it is challenging because the videos vary greatly in length and action instances occur more frequently in each video.
ActivityNet-v1.2(Caba Heilbron et al. 2015) covers 9682 videos that include 100 different classes of actions. These videos were categorized into 4619 training videos, 2383 validation videos, and 2480 test videos. Almost all of the videos contain at least one action category, and the action duration usually occupies more than half of the video duration. We performed model training on the training set and performance evaluation on the validation set.
Evaluation metrics. Following the standard protocol for temporal action localization [6], we chose the mean accuracy at different intersection-union ratio (IoU) thresholds (mAP@IoU) to evaluate the model performance. For a fair comparison, we use an IoU threshold range of [0.1:0.1:0.7] on the THUMOS-14 dataset and an IoU threshold range of [0.5:0.05:0.95] on ActivityNet-v1.2. The official evaluation code provided by ActivityNet is used to assess the model’s performance(Caba Heilbron et al. 2015).
@@
Comparison with State-of-the-art Methods.
In this section, we compare our proposed method with the current state-of-the-art temporal action localization methods, as presented in Tables 1 and 2. For the THUMOS14 dataset, we selected DELU and Zhou et al.(Zhou et al. 2023) as our baseline methods. After incorporating our MLLM4WTAL, we achieved improvements of 1.8% and 1.0% in the average mAP (0.1:0.7) metric, respectively, surpassing current weakly supervised methods and even outperforming some fully supervised methods, such as SSN, P-GCN, and G-TAD. On the larger-scale ActivityNet-v1.2 dataset, our method achieves a 0.5% improvement in average mAP compared to existing state-of-the-art methods.
Ablation Study
To analyze the contribution of each of our proposed modules, we conducted extensive ablation experiments to evaluate our approach from multiple perspectives, as shown in 3.
Baseline: In the baseline experimental setup, without using any of our proposed components, we achieved an average mAP of 37.4%.
Impact of KSM: In the Baseline+KSM experimental setup, where we solely incorporated the proposed Key Semantic Matching (KSM), we achieved an average mAP of 38.9%. The introduction of KSM resulted in a more significant mAP improvement, especially in high IoU settings. For instance, at IoU=0.7, there was a relative improvement of 1.6% compared to the Baseline.
Impact of CSR: In the Baseline+CSR experimental setup, where only the proposed Complete Semantic Reconstruction (CSR) was utilized, we achieved an average mAP of 38.7%. Similar to KSM, the introduction of CSR led to a more pronounced mAP improvement, with a relative increase of 2.2% at IoU=0.7 compared to the Baseline.
Combined Impact of KSM and CSR: In the Baseline+KSM+CSR experimental setup, where both KSM and CSR were simultaneously employed, we achieved an average mAP of 39.2%. The simultaneous introduction of KSM and CSR resulted in a more noticeable mAP improvement, particularly in high IoU settings. For example, at IoU=0.7, there was a relative improvement of 3.1% compared to the Baseline.
Generic validation

### v2

# Weakly Supervised Temporal Action Localization via Dual-Prior Collaborative Learning Guided by Multimodal Large Language Models

@@ Abstract
1 Introduction
Temporal action localization (TAL) [24] aims to localize action instances of interest from untrimmed videos. Although current fully-supervised TAL methods achieve significant localization results, fully-supervised methods require expensive, time-consuming, and labor-intensive frame-level annotations, and weakly-supervised temporal action localization (WTAL) methods [45], which require only video-level labels, have received much attention in order to alleviate the cost of annotation.
Existing WTAL methods typically train a classifier[45] using video-level labels, generating a set of probability predictions. This process results in temporal class activation maps (T-CAM). Despite notable performance improvements, current methods still face two challenges: incomplete and over-complete localization. As illustrated in Fig.1, some sub-actions with low discriminability may be ignored, leading to incomplete localization, while certain background segments that contribute to classification might be misclassified as foreground actions, causing over-complete localization.
@@ 1 Introduction
To overcome these challenges, we turn our perspective to existing multimodal large language models(MLLMs)[2]. The state-of-the-art MLLM employs a seamless amalgamation of a video foundation model and a large language model, culminating in the development of a sophisticated system for enhanced video comprehension. This integrated framework transcends the limitations inherent in traditional predefined visual tasks. Although MLLM exhibits powerful video understanding capabilities, the resource-intensive nature of running it poses high memory and computational costs. Therefore, traditional models remain essential for video understanding tasks. Is it possible to propose a learning paradigm that leverages MLLMs to enhance existing traditional WTAL methods? Inspired by this, we introduce a paradigm for MLLM-enhanced WTAL tasks, known as MLLM4WTAL.
@@
2 Related Work
Temporal Action Localizationl requires the simultaneous prediction of action start and end times along with the corresponding category. In fully-supervised temporal action localization training, frame-level labeling is commonly employed. Inspired by the widespread application of deep learning techniques[43, 44, 8, 34, 35, 33], existing methods mainly adopt a deep learning-based technical approach, which can be divided into the following two categories: top-down and bottom-up. Top-down methods [6] transform action localization into a time-dimensional detection task, inspired by image object detection. These methods generate action proposals, followed by classification and regression on temporal boundaries. They can draw on state-of-the-art techniques for Object Detection. On the other hand, bottom-up methods[36] generate frame-level predictions, followed by action localization using cleverly designed post-processing techniques. However, this fully supervised approach relies heavily on frame-by-frame annotation, which is both expensive and impractical when dealing with long videos.
Weakly-Supervised Temporal Action Localization has garnered increasing attention within the research community, particularly due to its ability to learn from video-level category labels during training, thus circumventing the need for precise temporal annotations[24]. The initial approach was UntrimmedNet [31], which first addressed this problem by employing a Multiple Instance Learning (MIL) framework that selects foreground segments and combines them into action segments. Subsequent improvement methods, such as STPN [22], enhanced UntrimmedNet by introducing a sparse loss to make the selected segments more sparse. CoLA [41] utilizes contrastive learning to distinguish between foreground and background segments. In addition, UGCT [37] proposed an online pseudo-label generation mechanism with uncertainty-aware learning capability for pseudo-label supervision of attention weights.
@@ 2 Related Work
Multimodal Large Language Models have emerged as a powerful paradigm for tackling a wide range of tasks spanning multiple modalities. LLM [4] has recently excelled in Natural Language Processing (NLP) tasks. To integrate with other modalities, researchers have worked to construct MLLM [2, 19]. Flamingo [2] integrates pre-trained visual and linguistic models to achieve state-of-the-art performance using a small number of samples. BLIP-2 [19] proposes a generic and efficient pre-training strategy that utilizes frozen pre-trained image encoders and large linguistic models for visual language pre-training. MiniGPT4 [49] implements the system by aligning frozen visual coders and LLMs, Vicuna [9], using only one projection layer. Otter [16] demonstrates improved instruction-following and context-learning capabilities. In the domain of video, ChatVideo [30] implements user interaction with LLMs using the tracklet as the fundamental video unit. As for Video-LLaMA [42], it utilizes the pre-trained models ImageBind [10] and LLaMA [29] on top of BLIP-2, which further facilitates cross-modal training for video.
3 Methods
@@
Overview. The proposed MLLM4WTAL framework, illustrated in Fig.2, enhances WTAL through two key aspects: Key Semantic Matching (KSM) and Complete Semantic Reconstruction (CSR), leveraging the semantic priors of MLLM. A dual priors interactive enhancement strategy is employed for mutual enhancement of these priors.
In section 3.1, MLLM first processes the raw video input to extract a temporal action key description. A key semantic query prompt is then constructed and input into the text encoder to obtain the key semantic query. RGB and flow video features $X_{rgb}$ and $X_{flow}$ are passed into the video embedding module to generate video feature embeddings. The KSM module matches the key semantic query with video segments, generating a contrastive query response that activates the most relevant video segment. Attention weights are applied to suppress background segments’ responses.
In section 3.2, MLLM inputs the raw video to generate a complete temporal action description and masks key action words in the sentence. An attention mechanism gathers fragments related to the masked words, and a linguistic reconstructor predicts the masked words.
Finally, section 3.3 utilizes a dual priors interactive enhancement strategy to foster collaboration between KSM and CSR, improving the accuracy and completeness of the localization results.
3.1 Key Semantic Matching
In this section, we will introduce key semantic matching (KSM) to comprehensively activate key temporal action intervals. Specifically, KSM consists of a key semantic generator, video embedding module, text embedding module, and semantic matching module.
Key Semantic Generator. In the key semantic generator module, the first step is to construct key semantic prompts for MLLM, guiding the MLLM to generate key descriptions for temporal actions. It is worth noting that when building the prompts, we have prior knowledge of the action categories present in the video. We introduce this action category before our prompt template. The prompt template $P_{\text{key}}$ format is as follows:
@@ 3.1 Key Semantic Matching
Video Embedding Module. Similar to other WTAL models, the video embedding module consists of two 1D convolutions followed by ReLU and Dropout layers. We adopt the same strategy as in [3] to fuse the RGB and flow features to obtain the input $\mathrm{F}\in\mathrm{R}^{\mathrm{T}\times 2048}$ of the video embedding module.
@@
The corresponding video feature embedding $\mathrm{F}_{\mathrm{e}}\in\mathrm{R}^{\mathrm{T}\times 2048}$ can then be obtained by $F_{e}=\Phi_{\text{emb}}(F)$, where $\Phi_{\text{emb}}$ represents the video embedding module. In addition, following previous works [11, 13], an attention mechanism is used to generate attention weight $\mathcal{A}_{KSM}\in\mathrm{R}^{\mathrm{T}\times 1}$ for each video clip.
@@
where $T_{\text{start}}$ represents the [START] token that was randomly initialized, $T_{\text{context}}$ is the learnable textual context of length $N_{context}$, and $T_{\text{key}}$ is the key informative textual feature embedded with GloVe [23]. Additionally, the $(C+1)th$ additional context class embedding is initialized with 0. Subsequently, a transformer encoder, denoted as $\Phi_{\text{trans}}$, functions as the text embedding module to generate text queries. Importantly, it should be noted that the key semantic query $F_{\text{query}}$ can be obtained through $F_{\text{query}}=\Phi_{\text{trans }}\left(T_{\text{query}}\right)$, where $F_{\text{query}}\in\mathrm{R}^{(\mathrm{C+1})\times 2048}$.
Semantic Matching Module. The semantic matching module aims to activate video segment features associated with key semantic queries. Specifically, we perform inner product operation on video embedding feature $X_{v}$ and text embedding feature $X_{t}$ to generate segment-level video-text similarity matrix $M\in R^{T\times(C+1)}$. In addition to this, following [11], we also employ the attention weight $\mathcal{A}_{KSM}$ to suppress the background segment responses. The segment-level matching result $\hat{M}\in R^{T\times(C+1)}$ for background suppression can be obtained by $\hat{M}=\mathcal{A}_{KSM}*M$, where ${}^{\prime}*^{\prime}$ denotes the element-wise multiplication. Following [24], we employ top-k multi-instance learning to compute the matching loss. Specifically, we compute the average of top-k similarities on the time dimension of the response of a particular text query category as the video-level video-text similarity. For the j-th action category, the video-level similarities $\mathrm{S}_{\mathrm{j}}$ and $\hat{S}_{\mathrm{j}}$ are generated from $M$ and $\hat{M}$, respectively.
@@
3.2 Complete Semantic Reconstruction
@@
To explore the complete temporal intervals of video that are semantically related to text, CSR includes an attention mechanism similar to 3.1. The attention weights $\mathcal{A}_{CSR}$ of the CSR module can be obtained through the following equation:
@@
Text Embedding Module. We mask the key action words of the video-complete description and embed the masked sentences with GloVe[25] and a fully connected layer to obtain the complete semantic embedding $\mathrm{F}_{\mathrm{c}}\in\mathrm{R}^{\mathrm{m}\times 512}$, where m is the length of the sentence.
Transformer Reconstructor. The Transformer reconstructor is used to reconstruct masked description sentences in the Complete Semantic Reconstruction module. First, Following [40], we randomly mask 1/3 of the words in the sentence as alternative description sentences, which may lead to a high probability of masking action labeling text. Then, the transformer’s encoder is utilized to obtain foreground video features $\mathrm{F}_{fg}\in\mathrm{R}^{\mathrm{T}\times 512}$
@@
$\mathbf{P}\left(w_{i}\mid\mathbf{F}_{\text{complete}},\mathbf{F}_{c[0:i-1]}\right)=\operatorname{softmax}\left(\Phi_{\mathrm{FC}}(\mathbf{H})\right)$ (16)
Where $\Phi_{\mathrm{FC}}$ is the fully connected layer .
@@
3.3 Dual Prior Interactive Distillation
Following PivoTAL[21], we utilized a prior-guided localization head. The loss function for the localization head is as follows:
@@ 4 Experiment
Full SSN[47] ICCV2017 66.0 59.4 51.9 41 29.8 - - 55.7 - -
P-GCN[39] CVPR2019 69.5 67.8 63.6 57.8 49.1 - - 61.6 - -
Weak CoLA[41] CVPR2021 66.2 59.5 51.5 41.9 32.2 22.0 13.1 50.3 32.1 40.9
DCC[18] CVPR2022 69.0 63.8 55.9 45.9 35.7 24.3 13.7 54.1 35.1 44.0
Li et al.[17] CVPR2023 71.1 65.0 56.2 47.8 39.3 27.5 15.2 55.9 37.2 46.0
Wang et al.[32] CVPR2023 73.0 68.2 60.0 47.9 37.1 24.4 12.7 57.2 36.4 46.2
Ren et al.[26] CVPR2023 71.8 67.5 58.9 49.0 40.0 27.1 15.1 57.4 38.0 47.1
Ju et al.[14] CVPR2023 73.5 68.8 61.5 53.8 42.0 29.4 16.8 59.9 40.7 49.4
PivoTAL[21] CVPR2023 74.1 69.6 61.7 52.1 42.8 30.6 16.7 60.1 40.8 49.7
AHLM[1] ICCV2023 75.1 68.9 60.2 48.9 38.3 26.8 14.7 58.3 37.8 47.6
CASE[20] ICCV2023 72.3 67.1 59.2 49.4 37.7 24.2 13.7 57.1 36.8 46.2
ISSF[38] AAAI2024 72.4 66.9 58.4 49.7 41.8 25.5 12.8 57.8 37.6 46.8
DELU[7] ECCV2022 71.5 66.2 56.5 47.7 40.5 27.2 15.3 56.5 37.4 46.4
@@
Zhou et al.[48] CVPR2023 74.0 69.4 60.7 51.8 42.7 26.2 13.1 59.7 38.9 48.3
MLLM4WTAL (Zhou et al.) - 74.3 69.8 61.8 52.3 43.0 30.8 16.6 60.2 40.9 49.8
@@
CoLA[41] CVPR2021 42.7 25.7 5.8 26.1
ACGNet[15] AAAI2022 41.8 26.0 5.9 26.1
DGCNN[27] MM2022 42.0 25.8 6.0 26.2
P-MIL[26] CVPR2023 44.2 26.1 5.3 26.5
ECM[46] TPAMI2023 41.0 24.9 6.5 25.5
Ren et al.[26] CVPR2023 44.2 26.1 5.3 25.5
Ju et al.[14] CVPR2023 48.3 29.3 6.1 29.6
CASE[20] ICCV2023 43.8 27.2 6.7 27.9
DELU[7] ECCV2022 44.2 26.7 5.4 26.8
@@
Table 3: Evaluation of each component of our method.(Unless otherwise specified, the THUMOS14 dataset is assumed by default.)
@@
mAP@IoU(%)
Method 0.3 0.4 0.5 0.6 0.7 AVG
Class name 56.7 47.9 40.4 27.6 15.9 37.7
key semantic 57.3 48.2 42.4 29.7 18.4 39.2 (+1.5)
a video of the [CLS] 56.4 48.0 40.5 27.5 15.8 37.6
complete semantic 57.3 48.2 42.4 29.7 18.4 39.2 (+1.6)
Table 5: Ablation study on prior information in MLLM
@@
THUMOS14[12] contains 413 untrimmed videos covering 20 different categories of actions. On average, each video contains about 15 unique instances of actions. We trained using 200 validation videos and evaluated on 213 test videos. Despite the small size of the dataset, it is challenging because the videos vary greatly in length and action instances occur more frequently in each video.
ActivityNet-v1.2[5] covers 9682 videos that include 100 different classes of actions. These videos were categorized into 4619 training videos, 2383 validation videos, and 2480 test videos. Almost all of the videos contain at least one action category, and the action duration usually occupies more than half of the video duration. We performed model training on the training set and performance evaluation on the validation set.
Evaluation metrics. Following the standard protocol for temporal action localization [6], we chose the mean accuracy at different intersection-union ratio (IoU) thresholds (mAP@IoU) to evaluate the model performance. For a fair comparison, we use an IoU threshold range of [0.1:0.1:0.7] on the THUMOS-14 dataset and an IoU threshold range of [0.5:0.05:0.95] on ActivityNet-v1.2. The official evaluation code provided by ActivityNet is used to assess the model’s performance.
@@
4.1 Comparison with State-of-the-art Methods.
In this section, we compare our proposed method with the current state-of-the-art temporal action localization methods, as presented in Tables 1 and 2. For the THUMOS14 dataset, we selected DELU and Zhou et al.[48] as our baseline methods. After incorporating our MLLM4WTAL, we achieved improvements of 1.8% and 1.0% in the average mAP (0.1:0.7) metric, respectively, surpassing current weakly supervised methods and even outperforming some fully supervised methods, such as SSN and P-GCN. On the larger-scale ActivityNet-v1.2 dataset, our method achieves a 0.5% improvement in average mAP compared to existing state-of-the-art methods.
4.2 Ablation Study
To analyze the contribution of each of our proposed modules, we conducted extensive ablation experiments to evaluate our approach from multiple perspectives, as shown in Table 3.
Baseline. In the baseline experimental setup, without using any of our proposed components, we achieved an average mAP of 37.4%.
Impact of KSM. In the Baseline+KSM experimental setup, where we solely incorporated the proposed Key Semantic Matching (KSM), we achieved an average mAP of 38.9%. The introduction of KSM resulted in a more significant mAP improvement, especially in high IoU settings. For instance, at IoU=0.7, there was a relative improvement of 1.6% compared to the Baseline.
Figure 3: Visualization of ground-truth and predictions.
Impact of CSR. In the Baseline+CSR experimental setup, where only the proposed Complete Semantic Reconstruction (CSR) was utilized, we achieved an average mAP of 38.7%. Similar to KSM, the introduction of CSR led to a more pronounced mAP improvement, with a relative increase of 2.2% at IoU=0.7 compared to the Baseline.
Combined Impact of KSM and CSR. In the Baseline+KSM+CSR experimental setup, where both KSM and CSR were simultaneously employed, we achieved an average mAP of 39.2%. The simultaneous introduction of KSM and CSR resulted in a more noticeable mAP improvement, particularly in high IoU settings. For example, at IoU=0.7, there was a relative improvement of 3.1% compared to the Baseline.
Impact of prior information. In Table 5, We analyzed the impact of MLLM from two perspectives: (1)The impact of key semantic priors provided by MLLM. We conducted an ablation study by replacing the key semantic priors provided by MLLM with the names of temporal action categories. As shown in Table 5, the average mAP improved by 1.5% when using key semantic priors.(2)The impact of complete semantic priors provided by MLLM. We conducted an ablation study by replacing the complete semantic priors provided by MLLM with fixed-template temporal action descriptions. As shown in Table 5, the average mAP improved by 1.6% when using complete semantic priors.
4.3 Generic validation